AnakKrakatau1883:

Graph g(x), where f(x) = 4x − 2 and g(x) = f(x + 1). a line labeled g of x that passes through points 0, negative 2 and 1, 2 a line labeled g of x that passes through points 0, negative 3 and 1, 1 a line labeled g of x that passes through points negative 1, negative 2 and 0, 2 a line labeled g of x that passes through points 0, negative 1 and 1, 3

Tranquility:

We are told g(x) = f(x+1) and we are given f(x) = 4x - 2 To find g(x), we have to replace the 'x' in f(x) with 'x+1' Essentially: g(x) = f(x+1) = 4(x+1) - 2 Can you simplify that and then graph it?